# Speedups in nonequilibrium thermal relaxation: Mpemba and related effects

### v1

@@ 1 Introduction
If the Mpemba effect is real, is it due to some peculiar feature of water, or is some general principle at play? As journalist Philip Ball memorably stated in 2006, “it is not clear whether the explanation [of the Mpemba effect] would be trivial or illuminating” [2]. Then, in 2017, two groups independently made a case for a general scenario. One proposed a “Markovian scenario where a system can start “farther” from equilibrium and yet reach equilibrium faster [3]. A second showed numerically and analytically that the effect could occur in models of granular fluids [4]. Strikingly, both works predicted the possibility of an inverse Mpemba effect, where a cold system heats up more quickly than a hot one. These developments were followed by an influential 2019 paper that predicted the possibility of a strong Mpemba effect, where the speed-up in relaxation is exponentially faster than the generic relaxation rate to equilibrium [5]. This latter prediction was soon after observed experimentally in a colloidal-particle system [6]. Around the same time, the Mpemba effect in quantum systems began to be discussed [7] and, in 2024, observed [8, 9, 10].
@@ 1.1 Historical background: from Aristotle to the present
I am surprised also that you should speak of noting down the points in my book which you regard as having been falsified by experience, for I venture to assure you that there are no such false points in it, because I made all the observations for myself, and especially the one you mention concerning warm water freezing faster than cold water. In the book I wrote, not “hot and cold”, but “water which has been heated over a fire for a long time freezes more quickly than other water”. In order to perform this experiment properly, the water must be allowed to cool down after it has boiled, until it has attained the same degree of coldness as water from a spring. After taking the temperature with a thermometer, you should take some water from a spring and pour equal quantities of the two sorts of water into identical vessels. But there are some experiments which only a few people are capable of performing, and when they are performed badly, the result is often the very opposite of what it ought to be.
@@
Both experiments by Black and Descartes compare samples of water that have identical initial conditions, with one sample boiled and cooled and the other not. This setup differs from the common scenario for the Mpemba effect, which compares samples of water that start at two different temperatures. Nonetheless, these early observations demonstrate a persistent “folklore” of anomalous behavior connected with the freezing of water. A more detailed discussion of this “pre-history” of the Mpemba effect (albeit excluding the papers by Barker and Black [21]) has been given by Jeng [22].
@@ 1.2 The cooling and freezing of water
If early accounts of anomalies in the cooling and freezing of water amount to a combination of folklore and rudimentary experiments spread diffusely over the centuries, the modern interest in such phenomena dates more precisely to 1969, when Erasto Mpemba and Denis Osborne published the paper “Cool?” [1]. Mpemba was a Tanzanian schoolboy (in Form 3, equivalent to 9th grade) in 1963, with a class assignment to make ice cream. The procedure was to boil a recipe, let it cool down to room temperature, and then put it in a freezer. Worried that another boy would fill up the freezer before his recipe had cooled, Mpemba put the hot mixture directly in the freezer. Unexpectedly, his hot mixture froze before the other boy’s. With admirable tenacity in the face of skeptical teachers and classmates, Mpemba sought to understand why his mixture had frozen first. Later, in high school, he convinced a visiting scientist, Denis Osborne, to investigate further.
@@
Figure 1: Time to start freezing as a function of initial sample temperature. This figure is reprinted from [1].
@@ 1.3 The zoology of Mpemba-like effects
    We distinguish between nonmonotonic relaxation effects in cooling processes, to which we refer as “direct Mpemba effects”, nonmonotonic relaxation in heating processes, which are “inverse Mpemba effects” and effects where the quenched parameter is not the temperature. We refer to the latter as “Mpemba-like effects”.
@@
    The final state of the system after the end of the relaxation process can be equilibrium, namely thermal distributions (as was the case in most experiments done in water, magnetic alloys, and other systems), nonequilibrium steady states (as in driven granular gases), or neither (as in spin glass systems). We refer to these effects as “equilibrium Mpemba”, “steady state Mpemba” and “glassy Mpemba”, respectively.
@@
    In various scenarios, different types of initial conditions can be discussed. In systems that relax to a final equilibrium state corresponding to a final temperature $T_{f}$, it is natural to consider the initial conditions as equilibrium distributions corresponding to different temperatures. Therefore, when the system relaxes to an equilibrium distribution, it is assumed that the initial distribution is also an equilibrium distribution. However, in several systems as driven granular gases the final state is not an equilibrium, but rather a nonequilibrium steady state. In this case, there are two different types of initial conditions: (i) A steady state corresponding to a different value of the “temperature” parameter in the system, while keeping all other parameters fixed. This is the analog of the equilibrium case, where the different initial conditions are identical in all aspects except for their initial temperature. (ii) A generic nonequilibrium distribution. In these cases, one has to be careful in dictating which initial condition is “closer” to the final state of the system.
@@
    In all related phenomena, there is some notion of nonmonotonic relaxation time, but the way this time is defined and measured is drastically different in the various suggested setups: in some cases, it is the time to the phase transition that dictates if the effect exists or not [57, 58]; in others, it is the time for the system to reach the new equilibrium / nonequilibrium steady state; and in yet other cases, it is the crossing in the relaxation curves of some other property that signals the distance from the final state, e.g., energy in spin glass [59]. Although it is plausible that there are some connections among these cases—for example, a cup of water that freezes first also reaches the freezer’s temperature faster—we distinguish among these various definitions. When the relaxation time is measured to a first or second-order phase transition, we refer to the effect as “Mpemba effect through a first / second-order phase transition”. When the nonmonotonic relaxation is measured in the long time limit, namely in the final stage of the relaxation, it is referred to as “final relaxation Mpemba effect”. A very strict definition was suggested in [60], where a system is defined to have a “thermomajorization Mpemba effect” if the crossing through relaxation happens for every monotone, namely for any quantity that always decays monotonically with the relaxation dynamics. Lastly, when the effect is declared based on the crossing of some observed quantity as effective temperature or average energy, we refer to it as a “crossing Mpemba effect” for the relevant property, e.g., in the spin glass, there is “Energy crossing Mpemba effect”. Note that, in principle, a system may have one type of Mpemba effect without the other, e.g., a final relaxation Mpemba effect without the existence of the effect through a phase transition that can occur in the system. The exact relations between these, e.g., whether crossing the phase transition first implies faster relaxation to the steady state, are yet unknown.
@@
    Following [5], we distinguish between the common case where the relaxation timescale is independent of the initial condition but the specific coefficient changes – the “weak Mpemba effect”, and cases where for the specific initial condition the relaxation timescale itself has a jump in its value, which is referred to as the “strong Mpemba effect”. We note that although this distinction was so far made only for final relaxation [5] and distance crossing [61] Mpemba effects, it is very plausible that a similar distinction exists also in phase transition Mpemba effects.
@@
    Recently, several types of nonmonotonic relaxation effects were measured in quantum systems. At least in some of these cases [8, 62, 63], the quantum nature of the system is crucial for the existence of the effect. When this is the case, we refer to it as a “Quantum Mpemba effect”.
@@
    Relaxation phenomena can occur due to interaction of the system with a thermal environment with which it exchanges heat (and possibly - particles) or in a closed system of many interacting particles, where the relaxation is towards an equilibrium distribution for the single particle distribution. When a nonmonotonic relaxation happens in a system that is coupled to a thermal environment, we refer to the effect as a “thermal Mpemba effect”, and if the system is a closed system, we refer to the effect as a “non-thermal Mpemba effect”.
The above different ways of categorizing experimental, analytical and numerical observations of the effect should be combined. For example, the first experimental observation of an inverse Mpemba effect in Ref. [64] should be classified as a “final relaxation inverse equilibrium thermal classical Mpemba effect”, and in fact, both the strong and the weak versions of the effects were measured in this case. Driven granular gases have both direct and inverse steady-state distance crossing classical thermal Mpemba effects, where the crossing is either in the effective temperature of the system or some other parameter of the velocity distribution. However, it is not known if a weak only version exists for the macroscopic mean-field anti-ferromagnetic Ising model [5] or for a large 2D Ising model [65], where we only know how to find the strong effect. Similarly, it is not known if the direct first-order phase transition equilibrium effect in Clathrate hydrate is a weak or a strong effect.
@@ 1.4 Notation
We use arrows to denote classical vectors, $\vec{v}$, Dirac notation with “bra”, $\bra{v}$ |, and “ket”, |$\ket{v}$, to denote wavefunctions, bold for matrices, A, and add a “hat”, $\hat{\textbf{L}}$, or introduce script notation, $\mathcal{L}$, for operators. The partition function, $\mathcal{Z}$, and the Hamiltonian, $\mathcal{H}$, are denoted by script notation. Typically, the probability that a system is in a state $i$ is denoted by $p_{i}$. The probability density is also denoted with small $p$ or $f$ (a more customary notation in kinetic theory, used in Sec. 4). From the context, we make it clear whether $p$ denotes the probability itself or the probability density. The Boltzmann distribution corresponding to a temperature $T$ is denoted by $\pi(T)$ or $\vec{\pi}(T)$. Unless otherwise specified, we set $k_{B}=1$, where $k_{B}$ is the Boltzmann constant. The Planck constant, $\hbar$, is likewise set to unity, $\hbar=1$. The inverse temperature in units of $k_{B}$ is denoted by $\beta\equiv 1/(k_{B}T)$. The $i$-th Ising spin is denoted by $\sigma_{i}$. In the granular section, $\vec{\sigma}$ is the standard notation for a vector connecting two centers of granular particles. Unless otherwise noted, the mass of a colloidal particle, $m$, and the damping coefficient $\gamma$ are set to unity. The reaction rate in Markovian jumps, $\Gamma$, is set to unity.
@@ 2 Markovian framework
From a modern perspective, the observation that hot water can freeze sooner than cold water, when all other conditions are identical, is consistent with similar experimental and numerical observations in other systems; see Sec. 6 for discussion of the relevant experiments and Sec. 7 for numerical observations. A natural approach is to address such nonmonotonic relaxation dynamics from a more general, abstract point of view. The framework of Markovian dynamics [66] represents the ideal nonequilibrium abstract setting to uncover the basic mechanisms that enable the emergence of these phenomena. The Markovian approach to nonequilibrium statistical physics has successfully described a wide variety of complex phenomena far from equilibrium, from the micro to the macro scale, with discrete or continuous degrees of freedom for single-body and many-body systems.
@@
In this section, we discuss a general framework for anomalous relaxation in both classical and quantum setups. First, we introduce a simple classical setup to describe thermal quenches: a system prepared in equilibrium at a given temperature $T_{0}$ is coupled to a thermal bath that has a different temperature $T_{b}$. This coupling initiates a relaxation process towards a new equilibrium at temperature $T_{b}$ that is driven by exchanges of heat with the thermal bath. We introduce a Markovian characterization of the time evolution after a thermal quench. The description is based on a Master Equation for systems with discrete degrees of freedom or on a Fokker-Planck equation for systems with continuous degrees of freedom. We then give criteria for the existence of a Mpemba effect. In contrast to the case of water, these criteria are based not on the phase transition time but rather on the properties of the last stage of the relaxation. Nevertheless, as we discuss in Sec. 3, the Markovian framework gives insight into the traditional definition of the Mpemba effect, too. Several examples highlighting the different phenomenology encountered in Markovian systems are discussed. Finally, we provide an overview of analogous effects observed in quantum systems, highlighting the strong analogies in the characterization that can be issued with Markovian open quantum systems, as well as parallels with non-thermal effects in closed quantum systems. Not all the quantum effects discussed in this section are Markovian, but since most are, we discuss them all in this section.
@@ 2.1 Classical systems
Markovian dynamics can describe both discrete degrees of freedom (DoF)—e.g., in spin systems, or continuous DoF—e.g., in a Brownian particle. In what follows, we discuss both cases. Each concept is first defined with continuous degrees of freedom and then is followed by the discrete case.
We consider a system with $n$ continuous DoF, which we identify with a vector $\vec{x}=\{x^{1},\dots,x^{n}\}$ in phase space (or configuration space) $\Omega$. Each DoF represents an internal continuous parameter of the system, such as the coordinates and momenta of a classical particle. A point $\vec{x}\in\Omega$ is also referred to as a microscopic configuration, or a microstate of the system. The Hamiltonian of the system, $\mathcal{H}(\vec{x})$, explicitly depends on the DoF, and it determines the energy of the system at each microscopic configuration. For a system with discrete DoF, e.g., a system of spins, the configuration space, the set of all possible states of the system, is discrete. We can then denote each microstate, namely each possible configuration of the system, as $x_{i}$, with $i=1,2,\ldots,N$. The energy of the system in microstate $x_{i}$ is $E_{i}$.
@@
A system that is in contact with a bath at temperature $T_{b}$ relaxes, after long enough time, to a thermal equilibrium with the thermal bath, regardless of the initial condition we had prepared the system in. The distribution corresponding to the thermal equilibrium is referred to as Boltzmann distribution. For continuous DoF it is given by
$\pi(\vec{x};T_{b})\propto e^{-\frac{1}{k_{B}T_{b}}\mathcal{H}(\vec{x})},$ (1)
where $k_{B}$ represents the Boltzmann constant (which, from here on, we set to one for convenience). The normalization of the equilibrium distribution, known as the partition function, is fixed by integrating over all the microscopic configurations as
@@
with
$\mathcal{Z}(T_{b})=\sum_{i}e^{-\frac{E_{i}}{T_{b}}}.$ (4)

### v3

@@ 1 Introduction
If the Mpemba effect is real, is it due to some peculiar feature of water, or is some general principle at play? As journalist Philip Ball memorably stated in 2006, “it is not clear whether the explanation [of the Mpemba effect] would be trivial or illuminating” [2]. Then, in 2017, two groups independently made a case for a general scenario. One proposed a “Markovian scenario where a system can start “farther” from equilibrium and yet reach equilibrium faster [3]. A second showed numerically and analytically that the effect could occur in models of granular fluids [4]. Strikingly, both works predicted the possibility of an inverse Mpemba effect, where a cold system heats up more quickly than a hot one. These developments were followed by an influential 2019 paper that predicted the possibility of a strong Mpemba effect, where the speedup in relaxation is exponentially faster than the generic relaxation rate to equilibrium [5]. This latter prediction was soon after observed experimentally in a colloidal-particle system [6]. Around the same time, the Mpemba effect in quantum systems began to be discussed [7] and, in 2024, observed [8, 9, 10].
@@ 1.1 Historical background: from Aristotle to the present
I am surprised also that you should speak of noting down the points in my book which you regard as having been falsified by experience, for I venture to assure you that there are no such false points in it, because I made all the observations for myself, and especially the one you mention concerning warm water freezing faster than cold water. In the book I wrote, not “hot and cold,” but “water which has been heated over a fire for a long time freezes more quickly than other water.” In order to perform this experiment properly, the water must be allowed to cool down after it has boiled, until it has attained the same degree of coldness as water from a spring. After taking the temperature with a thermometer, you should take some water from a spring and pour equal quantities of the two sorts of water into identical vessels. But there are some experiments which only a few people are capable of performing, and when they are performed badly, the result is often the very opposite of what it ought to be.
@@
Both experiments by Black and Descartes compare samples of water that have identical initial conditions, with one sample boiled and cooled and the other not. This setup differs from the common scenario for the Mpemba effect, which compares samples of water that start at two different temperatures. Nonetheless, these early observations demonstrate a persistent “folklore” of anomalous behavior connected with the freezing of water. A more detailed discussion of this “pre-history” of the Mpemba effect (although not including the papers by Barker and Black [21]) has been given by Jeng [22].
@@ 1.2 The cooling and freezing of water
If early accounts of anomalies in the cooling and freezing of water amount to a combination of folklore and rudimentary experiments spread diffusely over the centuries, the modern interest in such phenomena dates more precisely to 1969, when Erasto Mpemba and Denis Osborne published the paper “Cool?” [1]. Mpemba was a Tanzanian schoolboy (in Form 3, equivalent to 9th grade) in 1963, who wanted to make ice cream at school. The procedure was to boil a flavored mixture, let it cool down to room temperature, and then put it in a freezer. Worried that another boy would fill up the freezer before his recipe had cooled, Mpemba put the hot mixture directly in the freezer. Unexpectedly, his hot mixture froze before the other boy’s. With admirable tenacity in the face of skeptical teachers and classmates, Mpemba sought to understand why his mixture had frozen first. Later, in high school, he convinced a visiting scientist, Denis Osborne, to investigate further.
As Osborne recounts in his introduction to the paper,
The headmaster of Mkwawa High School invited me to speak to the students on Physics and National Development …. One student raised a laugh from his colleagues with a question I remember as: “If you take two beakers with equal volumes of water, one at 35 C and the other at 100 C, and put them into a refrigerator, the one that started at 100 C freezes first. Why?”
… I promised I would put the claim to the test of experiment ….
@@
Figure 1: Time to start freezing as a function of initial sample temperature. Source: Reprinted with permission from [1].
@@ 1.3 The zoology of Mpemba-like effects
    We distinguish between nonmonotonic relaxation effects in cooling processes, to which we refer as “direct Mpemba effects,” nonmonotonic relaxation in heating processes, which are “inverse Mpemba effects” and effects where the quenched parameter is not the temperature. We refer to the latter as “Mpemba-like effects.”
@@
    The final state of the system after the end of the relaxation process can be equilibrium, namely thermal distributions (as was the case in most experiments done in water, magnetic alloys, and other systems), nonequilibrium steady states (as in driven granular gases), or neither (as in spin glass systems). We refer to these effects as “equilibrium Mpemba,” “steady state Mpemba” and “glassy Mpemba,” respectively.
@@
    In various scenarios, different types of initial conditions can be discussed. In systems that relax to a final equilibrium state corresponding to a final temperature $T_{f}$, it is natural to consider the initial conditions as equilibrium distributions corresponding to different temperatures. Therefore, when the system relaxes to an equilibrium distribution, it is assumed that the initial distribution is also an equilibrium distribution. However, in several systems as driven granular gases the final state is not an equilibrium, but rather a nonequilibrium steady state. In this case, there are two different types of initial conditions: (i) A steady state corresponding to a different value of the “temperature” parameter in the system, while keeping all other parameters fixed. This is the analog of the equilibrium case, where the different initial conditions are identical in all aspects except for their initial temperature. (ii) A generic nonequilibrium distribution. In these cases, one has to be careful in determining which initial condition is “closer” to the final state of the system.
@@
    In all related phenomena, there is some notion of nonmonotonic relaxation time, but the way this time is defined and measured is drastically different in the various suggested setups: in some cases, it is the time to the phase transition that dictates if the effect exists or not [57, 58]; in others, it is the time for the system to reach the new equilibrium / nonequilibrium steady state; and in yet other cases, it is the crossing in the relaxation curves of some other property that signals the distance from the final state, e.g., energy in spin glass [59]. Although it is plausible that there are some connections among these cases—for example, a cup of water that freezes first also reaches the freezer’s temperature faster—we distinguish among these various definitions. When the relaxation time is measured to a first or second-order phase transition, we refer to the effect as “Mpemba effect through a first / second-order phase transition”. When the nonmonotonic relaxation is measured in the long time limit, namely in the final stage of the relaxation, it is referred to as “final relaxation Mpemba effect.” A very strict definition was suggested in [60], where a system is defined to have a “thermomajorization Mpemba effect” if the crossing through relaxation happens for every monotone, namely for any quantity that is assured to decay.
@@
    In all related phenomena, there is some notion of nonmonotonic relaxation time, but the way this time is defined and measured is drastically different in the various suggested setups: in some cases, it is the time to the phase transition that dictates if the effect exists or not [57, 58]; in others, it is the time for the system to reach the new equilibrium / nonequilibrium steady state; and in yet other cases, it is the crossing in the relaxation curves of some other property that signals the distance from the final state, e.g., energy in spin glass [59]. Although it is plausible that there are some connections among these cases—for example, a cup of water that freezes first also reaches the freezer’s temperature faster—we distinguish among these various definitions. When the relaxation time is measured to a first or second-order phase transition, we refer to the effect as “Mpemba effect through a first / second-order phase transition.” When the nonmonotonic relaxation is measured in the long time limit, namely in the final stage of the relaxation, it is referred to as “final relaxation Mpemba effect.” A very strict definition was suggested in [60], where a system is defined to have a “thermomajorization Mpemba effect” if the crossing through relaxation happens for every monotone, namely for any quantity that always decays monotonically with the relaxation dynamics. Lastly, when the effect is declared based on the crossing of some observed quantity as effective temperature or average energy, we refer to it as a “crossing Mpemba effect” for the relevant property, e.g., in the spin glass, there is an “Energy crossing Mpemba effect.” Note that, in principle, a system may have one type of Mpemba effect without the other, e.g., a final relaxation Mpemba effect without the existence of the effect through a phase transition that can occur in the system. The exact relations between these, e.g., whether crossing the phase transition first implies faster relaxation to the steady state, are yet unknown.
@@
    Following [5], we distinguish between the common case where the relaxation timescale is independent of the initial condition but the specific coefficient changes, the “weak Mpemba effect,” and cases where for the specific initial condition the relaxation timescale itself has a jump in its value, which is referred to as the “strong Mpemba effect.” We note that although this distinction was so far made only for final relaxation [5] and distance crossing [61] Mpemba effects, it is very plausible that a similar distinction exists also in phase-transition Mpemba effects.
@@
    Recently, several types of nonmonotonic relaxation effects were measured in quantum systems. At least in some of these cases [8, 62, 63], the quantum nature of the system is crucial for the existence of the effect. When this is the case, we refer to it as a “Quantum Mpemba effect.”
  8. 8.
    Relaxation phenomena can occur due to interaction of the system with a thermal environment with which it exchanges heat (and, possibly, particles) or in a closed system of many interacting particles, where the relaxation is towards an equilibrium distribution for the single particle distribution. When a nonmonotonic relaxation happens in a system that is coupled to a thermal environment, we refer to the effect as a “thermal Mpemba effect,” and if the system is a closed system, we refer to the effect as a “non-thermal Mpemba effect.”
The above different ways of categorizing experimental, analytical and numerical observations of the effect should be combined. For example, the first experimental observation of an inverse Mpemba effect in Ref. [64] should be classified as a “final relaxation inverse equilibrium thermal classical Mpemba effect,” and, in fact, both the strong and the weak versions of the effects were measured in this case. Driven granular gases have both direct and inverse steady-state distance crossing classical thermal Mpemba effects, where the crossing is either in the effective temperature of the system or some other parameter of the velocity distribution. However, it is not known if a weak only version exists for the macroscopic mean-field antiferromagnetic Ising model [5] or for a large 2D Ising model [65], where we only know how to find the strong effect. Similarly, it is not known if the direct first-order phase transition equilibrium effect in clathrate hydrates is a weak or a strong effect.
@@ 1.4 Notation
We use arrows to denote classical vectors, $\vec{v}$, Dirac notation with “bra”, $\bra{v}$ |, and “ket”, |$\ket{v}$, to denote wavefunctions, bold for matrices, A, and add a “hat”, $\hat{\textbf{L}}$, or introduce script notation, $\mathcal{L}$, for operators. The partition function, $\mathcal{Z}$, and the Hamiltonian, $\mathcal{H}$, are denoted by script notation. Typically, the probability that a system is in a state $i$ is denoted by $p_{i}$. The probability density is also denoted with small $p$ or $f$ (a more customary notation in kinetic theory, used in Sec. 4). From the context, we make it clear whether $p$ denotes the probability itself or the probability density. The Boltzmann distribution corresponding to a temperature $T$ is denoted by $\pi(T)$ or $\vec{\pi}(T)$. The inverse temperature is denoted by $\beta\equiv 1/(k_{B}T)$, where $k_{B}$ is the Boltzmann constant. The Planck constant, $\hbar$, is set to unity, $\hbar=1$. The $i$-th Ising spin is denoted by $\sigma_{i}$. In the granular section, $\vec{\sigma}$ is the standard notation for a vector connecting two centers of granular particles. Unless otherwise noted, the mass of a colloidal particle, $m$, and the damping coefficient $\gamma$ are set to unity. The reaction rate in Markovian jumps, $\Gamma$, is set to unity.
@@ 2 Markovian framework
From a modern perspective, the observation that hot water can freeze sooner than cold water, when all other conditions are identical, is consistent with similar experimental and numerical observations in other systems; see Sec. 6 for discussion of the relevant experiments and Sec. 7 for numerical observations. A natural approach is to address such nonmonotonic relaxation dynamics from a more general, abstract point of view. The framework of Markovian dynamics [66] represents the ideal nonequilibrium abstract setting to uncover the basic mechanisms that enable the emergence of these phenomena. The Markovian approach to nonequilibrium statistical physics has successfully described a wide variety of complex phenomena far from equilibrium, from the micro to the macro scale, with discrete or continuous degrees of freedom, for single-body and many-body systems.
@@
In this section, we discuss a general framework for anomalous relaxation in both classical and quantum setups. First, we introduce a simple classical setup to describe thermal quenches: a system prepared in equilibrium at a given temperature $T_{0}$ is coupled to a thermal bath that has a different temperature $T_{b}$. This coupling initiates a relaxation process towards a new equilibrium at temperature $T_{b}$ that is driven by exchanges of heat with the thermal bath. We introduce a Markovian characterization of the time evolution after a thermal quench. The description is based on a master equation for systems with discrete degrees of freedom or on a Fokker-Planck equation for systems with continuous degrees of freedom. We then give criteria for the existence of a Mpemba effect. In contrast to the case of water, these criteria are based not on the phase transition time but rather on the properties of the last stage of relaxation. Nevertheless, as we discuss in Sec. 3, the Markovian framework gives insight into the traditional definition of the Mpemba effect, too. Several examples highlighting the different phenomenology encountered in Markovian systems are discussed. Finally, we provide an overview of analogous effects observed in quantum systems, highlighting the strong analogies in the characterization that can be issued with Markovian open quantum systems, as well as parallels with non-thermal effects in closed quantum systems. Not all the quantum effects discussed in this section are Markovian, but since most are, we discuss them all in this section.
@@ 2.1 Classical systems
Markovian dynamics can describe both discrete degrees of freedom (DoF) (e.g., in spin systems) or continuous DoF (e.g., in a Brownian particle). In what follows, we discuss both cases. Each concept is first defined for continuous degrees of freedom and then for the discrete case.
We consider a system with $n$ continuous DoF, which we identify with a vector $\vec{x}=\{x^{1},\dots,x^{n}\}$ in phase space (or configuration space) $\Omega$. Each DoF represents an internal continuous parameter of the system, such as the coordinates and momenta of a classical particle. A point $\vec{x}\in\Omega$ is also referred to as a microscopic configuration, or a microstate of the system. The Hamiltonian of the system, $\mathcal{H}(\vec{x})$, explicitly depends on the DoF and determines the energy of the system at each microscopic configuration. For a system with discrete DoF, e.g., a system of spins, the configuration space, the set of all possible states of the system, is also discrete. We can then denote each microstate, namely each possible configuration of the system, as $x_{i}$, with $i=1,2,\ldots,N$. The energy of the system in microstate $x_{i}$ is $E_{i}$.
@@
A system that is in contact with a bath at temperature $T_{b}$ relaxes, after long-enough time, to a thermal equilibrium with the thermal bath, regardless of its initial condition. The distribution corresponding to the thermal equilibrium is referred to as the Boltzmann distribution. For continuous DoF, it is given by
$\pi(\vec{x};T_{b})\propto e^{-\frac{1}{k_{B}T_{b}}\mathcal{H}(\vec{x})}.$ (1)
The normalization of the equilibrium distribution, known as the partition function, is fixed by integrating over all the microscopic configurations as
@@
with partition sum
$\mathcal{Z}(T_{b})=\sum_{i}e^{-\frac{E_{i}}{k_{B}T_{b}}}.$ (4)